OpenStudy (anonymous):

Solve the equation, using the domain of x = {-1, 0, 2}. 2x + 3y = 7

OpenStudy (nurali):

first solve for any value of x. starting with 2x + 3y = 7 subtract 2x from both sides of the equation, then divide both sides of the equation by 3. you'll end up with: y = (-2x + 7)/3 now just evaluate for each specific value of x. when x=-1, y=(-2*-1 + 7)/3 = 9/3 = 3, so you have the point (-1, 3) when x=0, y=7/3, so (0, 7/3) when x=2, y=1, so (2, 1)